Solve the equation by completing the square: Solve the equation by completing the square: 4x^2-4x-9=0
Answer by rothauserc(4718)

4x^2-4x-9=0
rewrite equation
4x^2-4x = 9
divide both sides of = by 4
x^2 -x = 9/4
divide coefficient of x by 2, equals 1/2 and 1/2 squared = 1/4
x^2 -x +1/4 = 9/4 + 1/4
factor equation
(x-1/2)^ = 10/4
take square of both sides of =
(x -1/2) = + or - square root (10/4) = (+ or - square root(10)) / 2
x = 1/2 (+ or - square root(10)) / 2
x = 1/2 + (square root (10)) / 2
x = 1/2 - (square root (10)) / 2